Physics Question help! Thank you so much!

1. The are of the disk bounded by the circle is quadrupled, what happens to the net electric potential at the center of the circle?

Assume that the total charge is different from zero.
Answer #1

Electric potential at the centre of a charged disc, V = q/2πεo R where q is the charge on the disc and R is the radius of the disc.

Area of the disc, A = πR2 :

When the area of the disc is quadrupled, 4A/A = (R1/R)2 : 4 = (R1/R)2 :R1/R = 2; R1 = 2R

V = q/2πεo R implies that the electric potential V α 1/R

If V1 is the new potential then, V1/V = R/R1 = 1/2; V1 = V/2

When the area of a disc is quadrupled the electric potential become half the initial value of potential.